# WeCKD: Weakly-supervised Chained Distillation Network for Efficient Multimodal Medical Imaging

### v1

@@ Abstract
Knowledge distillation (KD) has traditionally relied on a static teacher-student framework, where a large, well-trained teacher transfers knowledge to a single student model. However, these approaches often suffer from knowledge degradation, inefficient supervision, and reliance on either a very strong teacher model or large labeled datasets, which limits their effectiveness in real-world, limited-data scenarios. To address these, we present the first-ever Weakly-supervised Chain-based KD network (WeCKD) that redefines knowledge transfer through a structured sequence of interconnected models. Unlike conventional KD, it forms a progressive distillation chain, where each model not only learns from its predecessor but also refines the knowledge before passing it forward. This structured knowledge transfer further enhances feature learning, reduces data dependency, and mitigates the limitations of one-step KD. Each model in the distillation chain is trained on only a fraction of the dataset and demonstrates that effective learning can be achieved with minimal supervision. Extensive evaluations across four otoscopic imaging datasets demonstrate that it not only matches but in many cases surpasses the performance of existing supervised methods. Experimental results on two other datasets further underscore its generalization across diverse medical imaging modalities, including microscopic and magnetic resonance imaging. Furthermore, our evaluations resulted in cumulative accuracy gains of up to +23% over a single backbone trained on the same limited data, which highlights its potential for real-world adoption.
@@ 3.2 Dataset
It is worth noting that some of the datasets initially contained a limited number of images. To enhance the framework’s generalization and prevent overfitting, data augmentation techniques were applied to get the mentioned number of samples; however, no preprocessing for data quality enhancement was employed to assess the model more accurately on raw data. Additionally, the datasets are referred to by their original names as collected in their respective source directories (see Table 6 for the full list of subclasses).
@@ 6 Conclusion
Acknowledgments
The authors declare no financial conflicts of interest that could have influenced this work.

### v2

@@ Abstract
Knowledge distillation (KD) has traditionally relied on a static teacher-student framework, where a large, well-trained teacher transfers knowledge to a single student model. However, these approaches often suffer from knowledge degradation, inefficient supervision, and reliance on either a very strong teacher model or large labeled datasets. To address these, we present the first-ever Weakly-supervised Chain-based KD network (WeCKD) that redefines knowledge transfer through a structured sequence of interconnected models. Unlike conventional KD, it forms a progressive distillation chain, where each model not only learns from its predecessor but also refines the knowledge before passing it forward. This structured knowledge transfer further enhances feature learning and addresses the limitations of one-step KD. Each model in the chain is trained on only a fraction of the dataset and shows that effective learning can be achieved with minimal supervision. Extensive evaluation on six imaging datasets across otoscopic, microscopic, and magnetic resonance imaging modalities shows that it generalizes and outperforms existing methods. Furthermore, the proposed distillation chain resulted in cumulative accuracy gains of up to +23% over a single backbone trained on the same limited data, which highlights its potential for real-world adoption.
@@ 3.2 Dataset
It is worth noting that some of the datasets initially contained a limited number of images. To enhance the framework’s generalization and prevent overfitting, data augmentation techniques were applied to get the mentioned number of samples; however, no preprocessing for data quality enhancement was employed to assess the model more accurately on raw data. Additionally, the datasets are referred to by their original names as collected in their respective source directories.
@@ 6 Conclusion
Declarations
Conflict of interests: The authors declare no financial conflicts of interest that could have influenced this work.
Ethics approval and consent to participate: No additional ethics approval or consent was required.
Data availability: The study uses six publicly available datasets: (1) Datos [42], (2) Eardrum [43], (3) OtoMatch [44], (4) EardrumDS [45], (5) Brain Tumor dataset [46], and (6) Lung and Colon Cancer dataset [47].